You throw a rock with an initial horizontal velocity and it hits the ground at x meters away from your feet. What if: (a) You throw the same stone with the same initial conditions but instead you do this on a another planet with a mass 4 times the mass of the earth. With what factor will x change? (b) You throw the same stone with the same initial conditions but instead you do this on a another planet with a radius 7 times the radius of the earth. With what factor will x change?
Answer #1

Let the height from which rock is thrown is h

g = \frac{GM}{R^2}

t = \sqrt{\frac{2h}{g}}

x = u_x t = u_x \sqrt {\frac{2h}{g}}

Now , when thrown on other planet .   

g' = \frac{GM'}{R'^2}

a) M' = 4M

g' = \frac{4GM}{R^2} = 4g

x' = u_x t = u_x \sqrt {\frac{2h}{g'}} = u_x \sqrt {\frac{2h}{4g}}

x' = \frac{x}{2}

b) . R' = 7R

g' = \frac{GM}{49R^2} = \frac{g}{49}

x' = u_x t = u_x \sqrt {\frac{2h}{g'}} = u_x \sqrt {\frac{98h}{g}}

x' = 7x